Find the measure of two complementary angles if one angle is 10 degress more than seven times the other. What are the measures of the two angles?
Answer by nerdybill(7384)

Find the measure of two complementary angles if one angle is 10 degress more than seven times the other. What are the measures of the two angles?
.
Let x = one angle
then
7x+10 = second angle
.
because they are complementary we get:
x + 7x+10 = 90
8x+10 = 90
8x = 80
x = 10 degrees
.
second angle:
7x+10 = 7(10)+10 = 70+10 = 80 degrees
.
answer: 10 and 80 degrees